Identify the center and radius of the circle with the given equations.
x^2 + y^2 + 8y -9 = 0

Answer by ikleyn(52803)

.

Transform given equation to the standard equation of a circle.


    x^2 + y^2 + 8y -9 = 0


    x^2 + y^2 + 8y = 9


    x^2 + (y^2 + 8y + 16) = 9 + 16


    x^2 + (y^2 + 4)^2 = 25


The circle with the center at (0,-4) and the radius of 5.    ANSWER

Solved.